OpenStudy (anonymous):

How would I solve this? :) Given: AB=x+7, BC=11, and BC=3x-8 (Drawing the pic) ^_____^

OpenStudy (anonymous):



OpenStudy (anonymous):

BC = 3x-8 , BC=11 ie 3x-8 = 11 x=(11+8)/3 = 19 / 3 = 6.33 sub x in AB then we get AB= 6.33 +7 = 13.33

OpenStudy (anonymous):

is this u expected??

OpenStudy (anonymous):

No, I wasn't sure how to do it. I was just doing whatever came to mind. I got it wrong but I was wondering what the process was to solve it. :) Thank you, though. ^____^

hero (hero):

Are you sure you posted your question correctly?

hero (hero):

You have two BC values. Are you solving for x or are you solving for AC?

OpenStudy (anonymous):

?!?!? Omg no it's wrong!!!!! I'm sorry!!! Thank you, Hero!!!! x(

OpenStudy (anonymous):

Given: AB=x+7, BC=11, and AC=3x-8

hero (hero):

:D

OpenStudy (anonymous):

Ok, I fixed it. :)

hero (hero):

So AB + BC = AC x+7 + 11 = 3x - 8 x + 18 = 3x - 8 2x = 26 x = 13

hero (hero):

There we go

OpenStudy (anonymous):

So, x is 13?

hero (hero):

Yup

OpenStudy (anonymous):

:D Ok, thank you both for helping me. ^____^